# Overcoming Label Noise for Source-free Unsupervised Video Domain Adaptation

### v1

@@ IV-D Hyperparameter Search
IV-E Impact of Teacher-Student Framework
IV-F Cross-domain Video Retrievals
In Table VI, we show the quantitative results for the cross-domain video retrieval task for the UCF $\leftrightarrow$ HMDB dataset. Our model retrieves better source videos from the target queries with R@1 of 0.92 and 0.91 as compared to the source-only model, which achieves only 0.82 and 0.88 on UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets respectively. In Figure 9, we show qualitative retrieval results for the UCF $\rightarrow$ HMDB. Our model can correctly retrieve the source videos of the same semantic categories as the target query videos.
As shown in Table VII, our proposed approach achieves better retrieval performance in most of the cases than the source-only model for the EPIC-Kitchen dataset. Only for D2 $\rightarrow$ D3, our model under-performs the source-only model. This can be attributed to the fact that our model does not use source data during the adaptation stage, and thus, the model might start forgetting some attributes of the source dataset.
V Conclusion
In this work, we address the relatively under-explored problem of source-free video domain adaptation and propose two simple yet effective approaches CleanAdapt and CleanAdapt + TS. Our framework is based on self-training in which we generate noisy pseudo-labels for the target domain data using the source pre-trained model. We argue that the presence of noise in the pseudo-labels hinders the adaptation performance and exploits the deep memorization effect to select clean samples in order to increase the quality of the pseudo-labels. Furthermore, we demonstrate that the teacher-student framework enhances the overall reliability of the adaptation training process, resulting in further performance improvements. Our methods consistently outperform the state-of-the-art image-based and video-based UDA methods without any source domain videos. We hope this perspective for video domain adaptation will help us approach other domain adaptation settings for videos.

### v2

@@ 4.6 Cross-domain Video Retrievals
As shown in Table 9, our proposed approach achieves better retrieval performance in most of the cases than the source-only model for the EPIC-Kitchen dataset. Only for D2 $\rightarrow$ D3, our model under-performs the source-only model. This can be attributed to the fact that our model does not use source data during the adaptation stage, and thus, the model might start forgetting some attributes of the source dataset.
4.7 Limitations and Future Work
In our model, we do not explicitly incorporate the spatiotemporal relationships inherent in videos. This information is crucial for capturing how objects and actions evolve over time, which can significantly enhance the model’s ability to understand complex video content. Without modeling these relationships, the model may struggle to effectively adapt to more challenging tasks such as video object segmentation, where it is essential to accurately track and delineate objects across frames.
A key assumption of our method is that each sample in both the source and target domains is associated with a single label and that they share a common label space $C$. While this is a standard setup in the literature, it may not always apply. For instance, CharadesEgo [sigurdsson2018charades] contains both first- and third-person videos, with each video having multiple labels. Our proposed approaches cannot be directly applied to such scenarios, and extending them to handle multi-label settings is left as future work.
However, we believe that the fundamental concept of selecting clean samples from noisy pseudo-labeled data will still be advantageous for tasks beyond video domain adaptation.
5 Conclusion
In this work, we address the relatively under-explored problem of source-free video domain adaptation and propose two simple yet effective approaches: CleanAdapt and CleanAdapt + TS. Our framework is based on self-training in which we generate noisy pseudo-labels for the target domain data using the source pre-trained model. Moreover, if we can filter out the noisy samples with varying $\tau$ using our proposed approach and use only the clean samples for fine-tuning, we achieve state-of-the-art performance without any video-specific modeling. To mitigate this issue of noisy pseudo-labels impeding the adaptation performance, we leverage the deep memorization effect [arpit2017closer] to identify and select the clean samples. Furthermore, we demonstrate that the quality of the clean samples can be improved by introducing a teacher-student framework, which in turn enhances the overall reliability of the adaptation training process and results in further performance improvements. Our methods consistently outperform recent image-based and video-based UDA methods without any source domain videos, thus establishing a new state of the art across several benchmarks.
[…]